# Collective action and entanglement of magnetically active liquid crystal elastomer ribbons

### v1

@@ 4 Experimental Section
Liquid crystal monomer 1,4-bis-[4-(6-acryloyloxyhexyloxy)benzoyloxy]-2-methylbenzene (RM82) was purchased from Henan Daken Chemical Co.,Ltd and monomer 1,4-bis-[4-(3-acryloyloxypropyloxy)benzoyloxy]-2-methylbenzene (RM257) was purchased from Wilshire inc. KG. Photoinitiators Irgacure I-369 and Irgacure I-651, chain extender n-butylamine, dimethylformamide, toluene, acetone, isopropanol, silicone oil (49 cP) were purchased from Fisher Scientific. Low viscosity silicone oil (8.8 cP) was purchased from Clearco inc. Brilliant yellow dye was purchased from Chem-Impex inc. Neodymium-iron-boron (NdFeB) magnetic microparticles (average diameter = 25 µm) were purchased from Magnequench. A two-part polydimethylsiloxane (PDMS) elastomer was purchased from Ellsworth Adhesives. High temperature block magnets (BX088SH) were purchased from k&J Magnetics.
@@
Clean glass slides (VWR) with dimensions of 38mm75mm1mm were subjected to O2 plasma for 5 min. Subsequently, the slides were spin-coated with 1wt% brilliant yellow solution in dimethylformamide at a spin speed of 1,500 RPM for 30 s.
The slides were baked for 30 min at 100°C on a hotplate to ensure the evaporation of dimethylformamide.
Each coated glass slide was exposed to linearly polarized broadband visible light from a modified projector at an intensity of 10 mW/cm2 for 7 min to align the brilliant yellow molecules [31].
For the preparation of polydomain samples, no polarized light exposure was used.
After alignment, the coated glass slides were spin-coated at a speed of 1,000 RPM for 15 s with a 9 wt% RM257 (reactive mesogens) and 1 wt% I-651 (photoinitiator) toluene solution (10 wt% solids).
The slides were then irradiated with 365 nm ultraviolet (UV) light at an intensity of 2 mW/cm–2 for 5 min using a UV oven (UVP Crosslinker CL-3000) for polymerization.
Glass cells were prepared by adhering a pair of RM257-coated glass slides together using a 50 µm Kapton Polyimide film spacer (McMaster Carr) and superglue (Gorilla).
A 1.4:1.0 molar ratio of RM82 to n-butylamine was heated and mixed with 1 wt% of I-369.
The solution was then allowed to partially fill the glass cell through capillary action on a hotplate at 75 °C before placing it in an oven at 75 °C to complete the filling and oligomerization.
After 15 h of oligomerization, the glass cells were allowed to cool to 40 ∘C temperature before exposing the cell to 365 nm UV light (Lumen Dynamics, OmniCure LX400+) at an intensity of 12 mW/cm–2 for polymerization for 50 min on each side.
After polymerization, an additional slide coated with brilliant yellow and polymerized RM257 (as mentioned above) was attached to the cell in the plane of one of the glass slides to create space for the magnetic domain.
The slide was separated using a 100 µm Kapton Polyimide film spacer (McMaster Carr) and adhered to the cell using superglue.
The formed space was filled with a 1.4:1.0 molar ratio of RM82 to n-butylamine with 20 wt% (NdFeB) magnetic microparticles and 1 wt% of I-369.
An additional oligomerization cycle was performed as mentioned above before exposure to 365 nm UV light (Lumen Dynamics, OmniCure LX400+) at 40 ∘C under an intensity of 12 mW/cm–2 for polymerization for 15 min on each side.
After polymerization, the top glass slide was released, and the film was cut using a CO2 laser cutter (Gravograph CO2 laser LS100-40W).
The cut ribbons were washed in water to remove brilliant yellow residue and the magnetic domains were magnetized under a 2.8 T impulse magnetic field (ACS Scientific IM-10-30).
The twisted nematic ribbons were magnetized at an angle corresponding to the offset angle with respect to the long axis of the ribbon.
Polydomain ribbons were magnetized in the direction parallel to the long axis of the ribbon.
@@
The apparatus frame was made from aluminium and constructed using a custom design.
A uniform magnetic field was achieved by forming a circular Halbach array (Halbach cylinder, ) made from an aluminium frame housing 16 Nickel-coated high temperature block magnets (N42SH) with dimensions 1"1/2"1/2" arranged in a circle with a 115 mm diameter, Figure S7.
The array was mounted on a 24V brushed DC Electric Motor.
The bottom side of a thin aluminium plate with dimensions 150 mm150 mm2 mm was mounted with a 6" adhesive silicone rubber heater 120 VAC, 10 W/in2 (35765K423, McMaster-Carr).
The plate was mounted 15 mm above the Halbach array, resulting in an average field strength of approximately 20 mT in the plane of plate.
The heater was connected to a Process PID controller (48VFL, EXTECH) powered by a variable AC voltage Transformer (TDGC-0.5KM, VEVOR) set to 94 V.
The oil temperature was checked using a thermocouple (20250-91, Digi-Sense).
@@
Curvature and degree of twist per unit length were measure by fixing a small (5 mm) piece of ribbon in place using a clip submerged in a temperature controlled oil bath.
The induced curvature at different temperatures was evaluated by fitting a circle to the curved surface of the actuated ribbon.
The degree of twist was evaluated by tracking the angle between the free end and the fixed end of the ribbon.
Three separate batches of ribbons were tested for each condition.
@@
Videos of experiments were captured using a Canon Rebel T5i camera from top view.
Tracking of the magnetic domain was conducted on 20 s videos taken at 60 FPS using a custom Matlab code.
The code captured the and coordinates of the average center of the magnetic domain in time.
To simplify the result and elucidate the features directly related to the magnetic field, the obtained time series were filtered using Matlab’s ‘Bandpass’ subroutine in the range [0.5,1.4] Hz to attenuate lateral motions that occur on different time scales than that of the field rotation, set to 1.2 Hz or 72 RPM.
Aggregation experiments were performed by capturing 2 min videos taken at 60 FPS. Analysis was applied on 40 frames taken at equal 3 s intervals by manual sampling of individual clusters and ribbons.
Dissociation experiments were performed by capturing 1 min videos taken at 60 FPS. Analysis was applied on 10 frames taken at equal 6 s intervals by manual sampling of individual clusters and ribbons.
To account for the effect of the viscosity, i.e., ’step-out’ frequency, the motion of a single ribbon in the field was tracked in time at different field rotation speeds.
Then, the power spectrum of the signal was obtained using the Matlab ‘pspectrum’ subroutine, and the largest non-zero frequency peak value was considered for the rotational frequency response of the ribbon.
Three separate batches of ribbons were tested for each condition.
@@
To measure the rheological properties of aggregates or dispersions, previously reported experiments were followed with modification [6]. Briefly, parallel-plate rheological scans were conducted utilizing a rheometer (Anton Paar Physica MCR 301). An aluminium well (50 mm diameter) was placed on the base to contain the ribbons and silicone oil in the rheometer (Figure 5). The oil temperature was checked using a thermocouple (20250-91, Digi-Sense). A consistent gap height of 3.5 mm was validated using the controlled rheometer displacement. Control experiments were performed in which no ribbons were placed in the oil (labeled ‘Only oil’) and in which the ribbons were dispersed in the way similar to the initial condition of the aggregation expeirments (labeled ‘No field’), Figure 5b. Aggregates were formed by dispersing the ribbons in a dish and applying the field rotation at 72 RPM for 4 minutes. Then, the formed aggregate was manually transferred using tweezers to the preheated aluminium well mounted on the rheometer.
Oscillatory strain and frequency sweeps with a gap of 3.5 mm were performed for measuring the storage and loss moduli as a function of shear strain from 0.01% to 100.00% (at an angular frequency of 10 rad/s–1), and 0.05 rad/s to 100 rad/s (at 0.05% strain).
Measurements taken at 90 ∘C were taken using a 3.35 mm gap height, approximately 4.5% lower than 3.5 mm, based on the average change in aggregate height associated with heating (Supplementary Figure S4).
Measurements taken at 120 ∘C were taken using a 3.18 mm gap height, approximately 9% lower than 3.5 mm, based on the average change in aggregate height associated with heating (Supplementary Figure S4).
The average changes in aggregate heights were obtained by measuring the dimension change relative to a submerged reference in a controlled oil heat bath.
The yield stress and yield strain values were measured as the shear stress and strain value at the cross-over point between the loss modulus and storage modulus [32, 33].
Three separate batches of ribbons were tested for each condition.
@@
Grazing-incidence wide-angle X-ray scattering (GIWAXS) was performed at the BioPACIFIC MIP user facilities at UC Santa Barbara.
GIWAXS was used to evaluate the alignment of LCE films.
Measurements were taken from homogeneously aligned (monodomain) nematic LCE films produced in the exact same way as described above, with the exception that the nematic director was uniform along the thickness of the film.
Grazing incidence experiments enable scattering measurements through the surface of a thin films, for which similar transmission geometry experiments would be cumbersome.
As the bottom half of the diffraction cone is absorbed by the substrate, only the upper half of the scattering pattern (Debye-Scherrer ring) is apparent (Supplementary Figure 1) [34].
The WAXS instrument was custom-built using a high brightness liquid metal jet X-ray source (Excillum MetalJet D2+ 70 keV) and a 4-megapixel hybrid photon-counting X-ray area detector (Dectris Eiger2 R 4M).
WAXS data were collected using a 9.24 keV beam in 10-minute intervals at an incidence angle of approximately 1∘ at room temperature.
Acknowledgements
@@
This work was supported National Science Foundation under Award No. DMR-2041671 and Texas A&M University. This work was partially supported by the BioPACIFIC Materials Innovation Platform of the National Science Foundation under Award No. DMR-1933487.
Conflict of Interest
The authors declare no conflict of interest.
Appendix A Supplementary information
A.1 List of supporting videos
Video S1: Entanglement and aggregation of ribbons.
Multi-angle view of aggregation of a ribbon dispersions with at C.
Video S2: Response to the magnetic field.
The trajectories of the magnetic domain of a single ribbon under different temperatures are tracked in time. The video shows the the tracked path (left panel), the actual motion of the ribbon (center panel), and the filtered path near 1.2 Hz (right panel) relating to the magnetic rotation.
Video S3: Aggregation experiment of different ribbon systems.
Comparison of representative aggregation experiments of: non-actuating ribbons (polydomain); at C; and at C.
Video S4: Simulations of dynamic entanglement
Represntative simulations of aggregation of ‘sticky’ and ‘non-sticky’ filaments with low ( at C) and high ( at C) curvatures.
Video S5: Aggregate response to cooling and heating.
Pre-formed aggregates ( at C) are cooled to C and heated to C.
Video S6: Aggregate dissociation.
Pre-formed aggregates ( at C) submerged in different viscosity oils are subjected to different field rotation speeds to induce dissociation.
A.2 Supplementary data
A.3 Mathematical model
We formulate a simplified mathematical model for the ribbon aggregation process by treating each ribbon as a rigid body subject to forces and torques. In particular, our model includes the effect of an external driving torque, arising from the magnetic field, a random force accounting for the drift of each ribbon, and contact forces and torques arising from their many-body interactions. The simulations shown in Figure 4 are performed by discretizing the ribbons, and then calculating the forces and torques on each ribbon in the discretized framework. Below we describe these forces and torques in more detail.
Treated as a rigid body, the state of a single ribbon can be constructed by rotating a translating a representative ribbon. Specifically, let for be the centerline of the ribbon, where is an arc length parameter in is the length of the ribbon. Suppose further that the curve is centered at the origin, so . The ’th ribbon at time , can then be described in terms of a rotation matrix and a translation
Setting , the ribbon velocity decomposes into rotational and translations parts
where the angular velocity can be written in index notation using the Levi-Civita symbol as . The angular velocity therefore satisfies
A state of the -ribbon system can be given in terms of rotations and translations
An allowed -ribbon configuration must also have the property that no two ribbons intersect each other. To capture contact, we approximate each ribbon as a tube with circular cross-sections of radius and with centerline . Using the approximation, an allowed configuration must satisfy the constraint
Additionally, the ribbons are in a domain . This constraint can be written
Each ribbon is subject to an external driving torque , an external stochastic drift force , as well as contact forces and torques arising from other ribbons and from the boundary . We treat the dynamics as overdamped and neglect hydrodynamics, obtaining the following simplified equations of motion
Although this framework corresponds to a simple diagonal resistance matrix, it is sufficient to capture the aggregation dynamics of the ribbons. The external forces and torque do not depend on the state of the ribbon aggregate, and therefore have relatively simple expressions. In particular, we take the external torque to be constant with magnitude
The external force is stochastic, given by
where are independent white noise processes satisfying
The fact that the -direction is treated differently from the -directions reflects the fact that the experiments typically take place in domains with horizontal extent much larger than vertical extent.
The expressions for contact forces and torques require a more careful accounting of the positions and overlaps of all the ribbons. There are several approaches for formulating these interactions [13, 27]. Here we construct contact forces by approximating each ribbon as a tube with circular cross-sections of radius (equation 3). We can quantify the contact between the ’th and ’th ribbons by setting and , and defining as follows
where is small. If , then when are touching and when overlap completely. Choosing has the desirable effect of smoothing out the contact region and the associated contact forces [28]. From we can construct a potential energy for soft contacts
where the parameter sets the magnitude of the repulsive contact forces and is a constant that sets the rate at which the potential stiffens. In addition, we assume that ribbons which are in contact must overcome an energy barrier in order to separate. This can be thought of as a ‘sticky’ adhesion force. A ribbon-ribbon contact which is on the point of separating has and . The additional adhesion term can therefore be taken to be proportional to where is the Heaviside step function. Including adhesion effects, the contact force density on the ’th ribbon due the ’th ribbon is then
where determines adhesion strength. From this expression we can write down the total contact force and torque on the ribbon
Finally boundary forces act to keep each ribbon confined to a domain . The boundary forces act at if or equivalent if . To write down the boundary force, for , let be the closest point to on the boundary of . Additionally, for , let be the inward normal to at . Then the boundary force density on the ’th ribbon is
where sets the strength of the boundary force. The boundary forces and torques are then
Our mathematical model for ribbon aggregation is given by equations (5)-(9).
A.3.1 Geometry and topology of ribbon aggregates
Geometric and topological quantities can be used to understand the structure of the ribbon aggregates formed by the model described above. For example the contact structure of these aggregates follows from constructing the contact graph, where each ribbon is a vertex and two vertices are joined by an edge if the corresponding ribbons are touching. More precisely, let be if ribbons and are touching, and otherwise. is therefore an adjacency matrix for the ribbon contact graph. The number of connected components of this graph provides a measure of aggregation of the ribbons.
The topological connectivity of ribbon aggregates can be quantified using the open linking number
For closed curves, the above integral yields a topological invariant whose value remains constant as the curves are deformed and moved, provided they do not pass through each other. For open curves, such as the ribbons we study here, this integral varies continuously as the open curves move. However, the open linking number nonetheless provides important information about how intertwined two curves are [13, 29]. Furthermore, the open linking numbers also allow us to measure the total amount of linking between the ’th ribbon and the rest of the aggregate
The average value of this parameter is the linking per filament, , which gives a measure of the overall total intertwining within an aggregate [13]
A random choice of ribbons within a domain will typically have small, nonzero linking per filament . This non-zero ‘background linking’ can be quantified finding the average value of for random collections of ribbons. Let be ribbons chosen uniformly at random in a domain , subject to the contact and boundary constraints described above (equation 3). The excess linking per filament can then be defined as
To avoid distortions due to the background linking, in Figure 4, we show the excess linking per filament.
A.4 Activation energy
The activation energy, , represents the minimal amount of rotational kinetic energy, , required to dissociate the physical bonds in the cluster. To estimate it, we employ the Arrhenius-like model
where [1/s] is the dissociation rate (measured number of steady state clusters per 20 s) and [1/s] is a constant that can be considered as the rate of collisions in the system (since in our system the rotational kinetic energy is the driving force for all types of lateral motions and subsequently collisions). We can approximate the rotational kinetic energy, , of a mildly coiled ribbon using the moment of inertia for a hoop rotating about its cylindrical axis, , where is the ribbon mass, is its radius, and is the magnetic field rotation speed. Using the formula for a circle with perimeter , the radius of the ribbon with mm can be estimated as approximately 2.5 mm (consistent with the curvature measurements in section 2.1). The linear regression (Figure 6) is then fitted to the Arrhenius plot for C yields the activation energy value nJ (95% CI: nJ), and 1/s. Since the critical activation energy required for a ribbon to induce dissociation also takes the form of a rotational kinetic energy, , we can substitute the fitted and previously calculated values to obtain RPM (95% CI: RPM), as heuristically estimated above. Similarly, the results for C yield nJ (95% CI: nJ) and 1/s, corresponding to RPM (95% CI: RPM).

### v2

@@ DISCUSSION
Liquid crystal monomer 1,4-bis-[4-(6-acryloyloxyhexyloxy)benzoyloxy]-2-methylbenzene (RM82) was purchased from Henan Daken Chemical Co.,Ltd and monomer 1,4-bis-[4-(3-
acryloyloxypropyloxy)benzoyloxy]-2-methylbenzene (RM257) was purchased from Wilshire inc. KG. Photoinitiators Irgacure I-369 and Irgacure I-651, chain extender n-butylamine, dimethylformamide, toluene, acetone, isopropanol, silicone oil ( mPas) were purchased from Fisher Scientific. Low viscosity silicone oil ( mPas) was purchased from Clearco inc. Brilliant yellow dye was purchased from Chem-Impex inc. Neodymium-iron-boron (NdFeB) magnetic microparticles (average diameter = 25 µm) were purchased from Magnequench. A two-part polydimethylsiloxane (PDMS) elastomer was purchased from Ellsworth Adhesives. High temperature block magnets (BX088SH) were purchased from k&J Magnetics. Glass microspheres (#22) were purchased from Fibre Glast.
@@
Clean glass slides (VWR) with dimensions of 38mm75mm1mm were subjected to O2 plasma for 5 min. Subsequently, the slides were spin-coated with 1wt% brilliant yellow solution in dimethylformamide at a spin speed of 1,500 RPM for 30 s. The slides were baked for 30 min at 100°C on a hotplate to ensure the evaporation of dimethylformamide. Each coated glass slide was exposed to linearly polarized broadband visible light from a modified projector at an intensity of 10 mW/cm2 for 7 min to align the brilliant yellow molecules 45. For the preparation of polydomain samples, no polarized light exposure was used. After alignment, the coated glass slides were spin-coated at a speed of 1,000 RPM for 15 s with a 9 wt% RM257 (reactive mesogens) and 1 wt% I-651 (photoinitiator) toluene solution (10 wt% solids). The slides were then irradiated with 365 nm ultraviolet (UV) light at an intensity of 2 mW/cm–2 for 5 min using a UV oven (UVP Crosslinker CL-3000) for polymerization. Glass cells were prepared by adhering a pair of RM257-coated glass slides together using a 50 µm Kapton Polyimide film spacer (McMaster Carr) and superglue (Gorilla). A 1.4:1.0 molar ratio of RM82 to n-butylamine was heated and mixed with 1 wt% of I-369. The solution was then allowed to partially fill the glass cell through capillary action on a hotplate at 75 °C before placing it in an oven at 75 °C to complete the filling and oligomerization. After 15 h of oligomerization, the glass cells were allowed to cool to 40 ∘C temperature before exposing the cell to 365 nm UV light (Lumen Dynamics, OmniCure LX400+) at an intensity of 12 mW/cm–2 for polymerization for 50 min on each side. After polymerization, an additional slide coated with brilliant yellow and polymerized RM257 (as mentioned above) was attached to the cell in the plane of one of the glass slides to create space for the magnetic domain. The slide was separated using a 100 µm Kapton Polyimide film spacer (McMaster Carr) and adhered to the cell using superglue. The formed space was filled with a 1.4:1.0 molar ratio of RM82 to n-butylamine with 20 wt% (NdFeB) magnetic microparticles and 1 wt% of I-369. An additional oligomerization cycle was performed as mentioned above before exposure to 365 nm UV light (Lumen Dynamics, OmniCure LX400+) at 40 ∘C under an intensity of 12 mW/cm–2 for polymerization for 15 min on each side. After polymerization, the top glass slide was released, and the film was cut using a CO2 laser cutter (Gravograph CO2 laser LS100-40W). The cut ribbons were washed in water to remove brilliant yellow residue and the magnetic domains were magnetized under a 2.8 T impulse magnetic field (ACS Scientific IM-10-30). The twisted nematic ribbons were magnetized at an angle corresponding to the offset angle with respect to the long axis of the ribbon. Polydomain ribbons were magnetized in the direction parallel to the long axis of the ribbon.
@@
The apparatus frame was made from aluminium and constructed using a custom design. A uniform magnetic field was achieved by forming a circular Halbach array (Halbach cylinder, ) made from an aluminium frame housing 16 Nickel-coated high temperature block magnets (N42SH) with dimensions 1”1/2”1/2” arranged in a circle with a 115 mm diameter, Figure S3. The array was mounted on a 24V brushed DC Electric Motor. The bottom side of a thin aluminum plate with dimensions 150 mm150 mm2 mm was mounted with a 6” adhesive silicone rubber heater 120 VAC, 10 W/in2 (35765K423, McMaster-Carr). The plate was mounted 20 mm above the Halbach array, resulting in an average field strength of approximately 26 mT in the plane of the plate, with good uniformity inside the area of a 60 mm ring (Figure S3). The heater was connected to a Process PID controller (48VFL, EXTECH) powered by a variable AC voltage Transformer (TDGC-0.5KM, VEVOR) set to 94 V. The oil temperature was checked using a thermocouple (20250-91, Digi-Sense).
Magnetic Characterization
The hysteresis loop of the magnetized LCE samples was measured with a vibrating sample magnetometer (VSM) (Quantum Design MPMS3) ranging from T to T. Three batches of composite LCE samples with 20wt% magnetic particles were synthesized and molded into flat films (using the method described above). The samples were folded to fit into the VSM sample holder. The magnetic moment of each sample was measured when the applied magnetic field was zero. The remanent magnetization of the composite was calculated by normalizing the measurement of magnetic moment to the sample weight ().
@@
Curvature and degree of twist per unit length were measure by fixing a small (5 mm) piece of ribbon in place using a clip submerged in a temperature controlled oil bath. The induced curvature at different temperatures was evaluated by fitting a circle to the curved surface of the actuated ribbon. The degree of twist was evaluated by tracking the angle between the free end and the fixed end of the ribbon. Three separate batches of ribbons were tested for each condition.
@@
Videos of experiments were captured using a Canon Rebel T5i camera from top view. Tracking of the magnetic domain was conducted on 20 s videos taken at 60 FPS using a custom Matlab code (Supplementary Information). The code captured the and coordinates of the average center of the magnetic domain in time. To simplify the result and elucidate the features directly related to the magnetic field, the obtained time series were filtered using Matlab’s ‘Bandpass’ subroutine in the range [0.5,1.4] Hz to attenuate lateral motions that occur on different time scales than that of the field rotation, set to 1.2 Hz (72 RPM). Aggregation experiments were performed by capturing 2 min videos taken at 60 FPS. Analysis was applied on 40 frames taken at equal 3 s intervals by manual sampling of individual clusters and ribbons. Dissociation experiments were performed by capturing 1 min videos taken at 60 FPS. Analysis was applied on 10 frames taken at equal 6 s intervals by manual sampling of individual clusters and ribbons. To account for the effect of the viscosity, i.e., ’step-out’ frequency, the motion of a single ribbon in the field was tracked in time at different field rotation speeds. Then, the power spectrum of the signal was obtained using the Matlab ‘pspectrum’ subroutine, and the largest non-zero frequency peak value was considered for the rotational frequency response of the ribbon. Three separate batches of ribbons were tested for each condition.
@@
To measure the rheological properties of aggregates or dispersions, previously reported experiments were followed with modification 6. Briefly, parallel-plate rheological scans were conducted utilizing a rheometer (Anton Paar Physica MCR 301). An aluminium well (50 mm diameter) was placed on the base to contain the ribbons and silicone oil in the rheometer (Figure 5). The oil temperature was checked using a thermocouple (20250-91, Digi-Sense). A consistent gap height of 3.5 mm was validated using the controlled rheometer displacement. Control experiments were performed in which no ribbons were placed in the oil (labeled ‘Only oil’) and in which the ribbons were dispersed in the way similar to the initial condition of the aggregation expeirments (labeled ‘No field’), Figure 5b. Aggregates were formed by dispersing the ribbons in a dish and applying the field rotation at 1.2 Hz (72 RPM) for 4 minutes. Then, the formed aggregate was manually transferred using tweezers to the preheated aluminum well mounted on the rheometer.
Oscillatory strain and frequency sweeps with a gap of 3.5 mm were performed for measuring the storage and loss moduli as a function of shear strain from 0.01% to 100.00% (at an angular frequency of 10 rad/s–1), and 0.05 rad/s to 100 rad/s (at 0.05% strain). Measurements taken at 90 ∘C were taken using a 3.35 mm gap height, approximately 4.5% lower than 3.5 mm, based on the average change in aggregate height associated with heating (Supplementary Figure S8). Measurements taken at 120 ∘C were taken using a 3.18 mm gap height, approximately 9% lower than 3.5 mm, based on the average change in aggregate height associated with heating (Supplementary Figure S8). The average changes in aggregate heights were obtained by measuring the dimension change relative to a submerged reference in a controlled oil heat bath. The yield stress and yield strain values were measured as the shear stress and strain value at the cross-over point between the loss modulus and storage modulus 46, 47. Three separate batches of ribbons were tested for each condition.
@@
Grazing-incidence wide-angle X-ray scattering (GIWAXS) was performed at the BioPACIFIC MIP user facilities at UC Santa Barbara. GIWAXS was used to evaluate the alignment of LCE films. Measurements were taken from homogeneously aligned (monodomain) nematic LCE films produced in the exact same way as described above, with the exception that the nematic director was uniform along the thickness of the film. Grazing incidence experiments enable scattering measurements through the surface of a thin films, for which similar transmission geometry experiments would be cumbersome. As the bottom half of the diffraction cone is absorbed by the substrate, only the upper half of the scattering pattern (Debye-Scherrer ring) is apparent (Supplementary Figure S1) 48. The WAXS instrument was custom-built using a high brightness liquid metal jet X-ray source (Excillum MetalJet D2+ 70 keV) and a 4-megapixel hybrid photon-counting X-ray area detector (Dectris Eiger2 R 4M). WAXS data were collected using a 9.24 keV beam in 10-minute intervals at an incidence angle of approximately 1∘ at room temperature.
Resource availability
Lead contact
Requests for further information and resources should be directed to and will be fulfilled by the lead contact, Taylor H. Ware (taylor.ware@tamu.edu).
Materials availability
This study did not generate new materials.
Data and code availability
The data reported in this paper is available from the lead contact upon request.
All original code relating to the mathematical model is publicly available at GitHub (github.com/vppatil28/rigid_filaments) as of the date of publication. All original code related to tracking and analysis of experimental data is attached in the supplementary information of this paper.
The Matlab code used to model the Halbach array 49 is available at
https://www.blogs.uni-mainz.de/fb08-physics-halbach-magnets/software-2/. -
Any additional information required to reanalyze the data reported in this paper is available from the lead contact upon request.
Supplemental information index
Videos S1-S9 in mp4 format.
Videos S1-S9 descriptions and their legends in a PDF.
Figures S1-S11 and their legends in a PDF.
Details on mathematical models in a PDF.
Original code used in this work in a PDF.
Acknowledgments
@@
We thank Justin M. Gonzalez and the Fischer Engineering Design Center team for their assistance with samples preparations. This work was supported National Science Foundation under Awards No. DMR-2041671 & CMMI-2427025 and Texas A&M University. This work was partially supported by the BioPACIFIC Materials Innovation Platform of the National Science Foundation under Award No. DMR-1933487.
Author contributions
Conceptualization, A.D., T.H.W., and M.K.A.; methodology, A.D., V.P.P., M.S.K., S.T.; investigation, A.D., C.B., K.H., D.C.C., J.A.G., P.A.K., Y.L., and V.P.P.; visualization, A.D., S.M.G., and M.S.K; writing – original draft, A.D. and T.H.W.; writing – review & editing, A.D., C.B., M.S.K., K.H., S.M.G., D.C.C., J.A.G., S.T., P.A.K., Y.L., M.K.A., V.P.P., and T.H.W.; funding acquisition, T.H.W.; resources, T.H.W, and A.D.; supervision, T.H.W.;
Declaration of interests
The authors declare no competing interests.